# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v1

@@ Appendix A Technical Lemmas
where the remainder $\varepsilon(\alpha,m,\eta)$ is the same as in Theorem 3.1, in particular $\varepsilon(\alpha,m,\eta)=O\big(\sqrt{\log m}\big)$ and $\varepsilon(\alpha,m,\eta)/\sqrt{m}=O\big(\sqrt{(\log m)/m}\big)$.
@@
The argument is identical to the proof of Theorem 3.1, except that we now construct separate confidence sets for $p_{j}$ and $q_{j}$ with marginal coverages $\gamma_{p}$ and $\gamma_{q}$ chosen so that $\gamma_{p}\gamma_{q}=\tfrac{1}{2}$. For example, we can set $\gamma_{p}=\gamma_{q}=\tfrac{1}{\sqrt{2}}$. This ensures that the joint event $\{p_{j}\in C_{j}^{p}(\hat{p}_{j}),\,q_{j}\in C_{j}^{q}(\hat{q}_{j})\}$ plays exactly the same role as the univariate coverage event in Theorem 3.1, i.e., $\mathbb{P}(\hat{\Delta}_{j}^{\star,(k)}\geq\Delta_{j}^{\star,(k)}\mid\mathcal{G}_{j})\geq\tfrac{1}{2}$. All subsequent steps then carry over verbatim, yielding the stated bound. ∎
B.3 Proof to Theorem 5.1:
@@
We follow the setup of Theorem 3.1. Recall that we work conditionally on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j}),\mathcal{G}:=\sigma\big(\{\mathcal{G}_{j}\}_{j=1}^{m}\big).$ For brevity, write $p_{j}:=p(\psi_{j})$, $q_{j}:=q(\psi_{j})$, and similarly $\hat{p}_{j},\hat{q}_{j}$. The true per-scenario discrepancy is $\Delta_{j}:=L(p_{j},\hat{q}_{j}),\text{ }j=1,\dots,m,$ and by the i.i.d. scenario assumption, $\{\Delta_{j}\}_{j=1}^{m}$ are i.i.d. draws from the distribution of $\Delta(\psi)$ under $\psi\sim\Psi$.
@@
By Assumption 3.2 and compactness of $C_{j}(\hat{p}_{j})$, Berge’s maximum theorem implies that the infimum is attained on the confidence set, hence $\Delta_{j}^{-}$ is well defined.
@@
Also by definition, on the event $\{p_{j}\in C_{j}(\hat{p}_{j})\}$ we clearly have $\Delta_{j}^{-}\leq\Delta_{j}.$ Thus, for each $j$,
$\mathbb{P}\big(\Delta_{j}^{-}\leq\Delta_{j}\big|\mathcal{G}_{j}\big)=\mathbb{P}\big(p_{j}\in C_{j}(\hat{p}_{j})\big|\mathcal{G}_{j}\big)\geq\tfrac{1}{2}.$
@@
Define the indicators $Y_{j}^{-}:=\mathbf{1}\{\Delta_{j}^{-}\leq\Delta_{j}\},\qquad j=1,\dots,m.$ Conditionally on $\mathcal{G}$, the variables $\{Y_{j}^{-}\}_{j=1}^{m}$ are independent and satisfy
$\mathbb{P}(Y_{j}^{-}=1\mid\mathcal{G})\ \geq\ \tfrac{1}{2},\qquad\forall j.$ (B.6)
@@
Fix $\alpha\in(0,1)$ for the moment. We again denote $\bar{V}_{m}(\alpha)$ as the empirical $\alpha$-quantile of the (unobservable) $\{\Delta_{j}\}_{j=1}^{m}$, i.e., $\bar{V}_{m}(\alpha):=\Delta_{(\lceil m\alpha\rceil)},$ where $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}$ are the order statistics, and let $\bar{F}_{m}(t):=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}\leq t\}$ be their empirical CDF.
Define the index set of “lower-tail” scenarios
@@
By definition of $\bar{V}_{m}(\alpha)$ we have $s^{-}\geq m\alpha$, and these are the true $\Delta$ quantile that we wish to capture. Among these, consider those for which the lower pseudo-gap is valid:
$T_{\alpha}^{-}$ $:=\{j\in S_{\alpha}^{-}:Y_{j}^{-}=1\}$
$=\{j\in S_{\alpha}^{-}:\Delta_{j}^{-}\leq\Delta_{j}\leq\bar{V}_{m}(\alpha)\}.$
By (B.6), $\mathbb{E}\big[|T_{\alpha}^{-}|\big|\mathcal{G}\big]=\sum_{j\in S_{\alpha}^{-}}\mathbb{P}(Y_{j}^{-}=1\mid\mathcal{G})\geq\tfrac{1}{2}s^{-}.$
Applying Lemma A.2, for any $t\in[0,s^{-}/2]$,
$\mathbb{P}\Big(|T_{\alpha}^{-}|\leq\tfrac{1}{2}s^{-}-t\Big|\mathcal{G}\Big)\leq\exp\Big(-\frac{2t^{2}}{s^{-}}\Big).$
@@
$\mathbb{P}\Big(|T_{\alpha}^{-}|\leq\tfrac{1}{2}s^{-}-c\sqrt{s^{-}}\Big|\mathcal{G}\Big)\leq e^{-2c^{2}}.$ (B.7)
@@
$|U_{\alpha}^{-}|\;\geq\;|T_{\alpha}^{-}|\;\geq\;\tfrac{1}{2}s^{-}-c\sqrt{s^{-}}\ \geq\ \tfrac{1}{2}m\alpha-c\sqrt{m\alpha}.$
Let $\hat{F}_{m}^{-}(t)$ denote the empirical CDF of $\{\Delta_{j}^{-}\}_{j=1}^{m}$, ie. $\hat{F}_{m}^{-}(t):=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}^{-}\leq t\},$ and let $\hat{V}_{m}^{-}(\beta)$ be its empirical $\beta$-quantile.
From the bound on $|U_{\alpha}^{-}|$ we obtain
$\hat{F}_{m}^{-}\big(\bar{V}_{m}(\alpha)\big)=\frac{|U_{\alpha}^{-}|}{m}\geq\tfrac{1}{2}\alpha-c\sqrt{\tfrac{\alpha}{m}}-\tfrac{1}{m}.$ (B.8)
Thus, for any
$\beta\leq\beta_{\mathrm{eff}}^{-}(\alpha,c,m):=\tfrac{1}{2}\alpha-c\sqrt{\tfrac{\alpha}{m}}-\tfrac{1}{m},$
we have $\hat{F}_{m}^{-}(\bar{V}_{m}(\alpha))\geq\beta$, and hence, by the definition of the empirical quantile,
$\hat{V}_{m}^{-}(\beta)\;\leq\;\bar{V}_{m}(\alpha),\qquad\forall\,\beta\leq\beta_{\mathrm{eff}}^{-}(\alpha,c,m).$ (B.9)
Let $\mathcal{E}_{\alpha}^{-}$ denote the event that both (B.7) (complement) and (B.9) hold; then $\mathbb{P}(\mathcal{E}_{\alpha}^{-}\mid\mathcal{G})\geq 1-e^{-2c^{2}}.$
Finally, we try to bridge $\bar{F}_{m}(t)$ to the true CDF of $\Delta(\psi)$ under $\psi\sim\Psi$, which we denote as $F_{\Delta}$. By DKW inequality (Lemma A.1), for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have the following event:
@@
$\varepsilon_{m}(\delta):=\sqrt{\frac{\log(2/\delta)}{2m}}.$
Work on the intersection $\mathcal{E}_{\mathrm{DKW}}\cap\mathcal{E}_{\alpha}^{-}$. Fix $\beta\leq\beta_{\mathrm{eff}}^{-}(\alpha,c,m)$ and set $t^{-}:=\hat{V}_{m}^{-}(\beta)$. By (B.9), we have $t^{-}\leq\bar{V}_{m}(\alpha).$ Using this and the fact that $F_{\Delta}$ is nondecreasing, we have
$F_{\Delta}(t^{-})\leq F_{\Delta}\big(\bar{V}_{m}(\alpha)\big).$
On $\mathcal{E}_{\mathrm{DKW}}$, the DKW bound yields
$F_{\Delta}\big(\bar{V}_{m}(\alpha)\big)\leq\bar{F}_{m}\big(\bar{V}_{m}(\alpha)\big)+\varepsilon_{m}(\delta)\leq\alpha+\varepsilon_{m}(\delta).$
Combining the last two equations,
$F_{\Delta}(t^{-})\leq\alpha+\varepsilon_{m}(\delta).$
@@
$F_{\Delta}\big(\hat{V}_{m}^{-}(\beta)\big)=\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\leq\hat{V}_{m}^{-}(\beta)\big|\mathcal{D}\big)\leq\alpha+\varepsilon_{m}(\delta),$
hence on the event $\mathcal{E}_{\mathrm{DKW}}\cap\mathcal{E}_{\alpha}^{-}$ we have,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\geq\hat{V}_{m}^{-}(\beta)\big|\mathcal{D}\big)$ $\geq 1-\alpha-\varepsilon_{m}(\delta),$ (B.10)
$\qquad\forall\beta\leq\beta_{\mathrm{eff}}^{-}(\alpha,c,m).$
Finally, we specify the actual adjustment term $\beta$ for a fixed $\alpha$. Fix $\alpha\in(0,1)$. We wish to apply (B.10) with $\beta=\alpha/2$ to keep the argument of $\hat{V}_{m}^{-}$ simple. Recall that (B.10) holds whenever $\beta\leq\beta_{\mathrm{eff}}^{-}(\alpha^{\prime},c,m)$ for some $\alpha^{\prime}$. Thus we need to choose $\alpha^{\prime}$ such that
$\tfrac{1}{2}\alpha^{\prime}-c\sqrt{\tfrac{\alpha^{\prime}}{m}}-\tfrac{1}{m}\geq\tfrac{\alpha}{2}.$
Writing $y:=\sqrt{\alpha^{\prime}}$, we obtain the quadratic inequality $y^{2}-\frac{2c}{\sqrt{m}}y-\Big(\alpha+\tfrac{2}{m}\Big)\geq 0.$ Solving the minimal admissible solution we get:
$y^{*}(\alpha,c,m)$ $:=\frac{\tfrac{2c}{\sqrt{m}}+\sqrt{\tfrac{4c^{2}}{m}+4\alpha+\tfrac{8}{m}}}{2},$
$\alpha_{\mathrm{eff}}^{-}(\alpha,c,m)$ $:=\big(y^{*}(\alpha,c,m)\big)^{2}.$
For this choice of $\alpha^{\prime}$ we have $\alpha/2\leq\beta_{\mathrm{eff}}^{-}(\alpha^{\prime},c,m)$, so (B.10) with $\beta=\alpha/2$ gives
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\geq\hat{V}_{m}^{-}(\tfrac{\alpha}{2})\,\big|\mathcal{D}\big)\geq 1-\alpha_{\mathrm{eff}}^{-}(\alpha,c,m)-\varepsilon_{m}(\delta).$
Explicitly solving quadratic solution, one can derive
$\alpha_{\mathrm{eff}}^{-}(\alpha,c,m)-\alpha\leq\frac{c}{\sqrt{m}}\sqrt{4\alpha+\tfrac{4c^{2}+8}{m}}+\frac{2c^{2}+2}{m},$
hence implies
$\mathbb{P}_{\psi\sim\Psi}\big($ $\Delta(\psi)\geq\hat{V}_{m}^{-}(\tfrac{\alpha}{2})\big|\mathcal{D}\big)\geq$
$1-\alpha-\frac{c}{\sqrt{m}}\sqrt{4\alpha+\tfrac{4c^{2}+8}{m}}-\frac{2c^{2}+2}{m}-\varepsilon_{m}(\delta).$ (B.11)
Again, we now make the bound uniform over $\alpha\in(0,1)$ and tie $(c,\delta)$ to a target failure probability $\eta\in(0,1)$.
Consider the grid $\alpha_{r}:=r/m$ for $r=1,\dots,m$. For each $r$, the fixed-level argument above yields an event
@@
From the Chernoff step we have, for each $r$, $\mathbb{P}(\mathcal{E}_{r}^{-})\ \geq\ 1-e^{-2c^{2}}.$ Hence, by a union bound,

### v2

@@ B.3 Corollary for simulator budget:
where the remainder $\varepsilon(\alpha,m,\eta)$ is the same as in Theorem B.1, in particular $\varepsilon(\alpha,m,\eta)=O\big(\sqrt{\log m}\big)$ and $\varepsilon(\alpha,m,\eta)/\sqrt{m}=O\big(\sqrt{(\log m)/m}\big)$.
@@
The argument is identical to the proof of Theorem B.1, except that we now construct separate confidence sets for $p_{j}$ and $q_{j}$ with marginal coverages $\gamma_{p}$ and $\gamma_{q}$ chosen so that $\gamma_{p}\gamma_{q}=\tfrac{1}{2}$. For example, we can set $\gamma_{p}=\gamma_{q}=\tfrac{1}{\sqrt{2}}$. This ensures that the joint event $\{p_{j}\in C_{j}^{p}(\hat{p}_{j}),\,q_{j}\in C_{j}^{q}(\hat{q}_{j})\}$ plays exactly the same role as the univariate coverage event in Theorem 3.1, i.e., $\mathbb{P}(\hat{\Delta}_{j}^{\star,(k)}\geq\Delta_{j}^{\star,(k)}\mid\mathcal{G}_{j})\geq\tfrac{1}{2}$. All subsequent steps then carry over verbatim, yielding the stated bound. ∎
Theorem B.1.
Suppose the setup of Theorem 3.1 and Assumptions 3.13.2 hold. Let $\{\gamma_{L,j},\gamma_{U,j}\}_{j=1}^{m}\subset(0,1]$ be coverage levels and define the lower and upper pseudo-discrepancies
$\hat{\Delta}_{j}^{-}:=\inf_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{L,j})}L(u,\hat{q}_{j}),\ \hat{\Delta}_{j}^{+}:=\sup_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{U,j})}L(u,\hat{q}_{j}),$
where $\mathcal{C}_{j}(\hat{p}_{j},\gamma)\subset\Theta$ are data-driven compact confidence sets satisfying the conditional coverage guarantees
$\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{L,j})\mid\psi_{j},n_{j}\big)$ $\geq\gamma_{L,j}$
$\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{U,j})\mid\psi_{j},n_{j}\big)$ $\geq\gamma_{U,j}.$
Let $\hat{V}_{m}^{-}(\cdot)$ denote the empirical quantile function of $\{\hat{\Delta}_{j}^{-}\}_{j=1}^{m}$, and $\hat{V}_{m}^{+}(\cdot)$ the empirical quantile function of $\{\hat{\Delta}_{j}^{+}\}_{j=1}^{m}$. Define the empirical mean coverage levels $\bar{\gamma}_{L,m}:=\frac{1}{m}\sum_{j=1}^{m}\gamma_{L,j},\bar{\gamma}_{U,m}:=\frac{1}{m}\sum_{j=1}^{m}\gamma_{U,j}.$ For $\alpha\in(0,1)$ and $\delta\in(0,1)$ define the discretized level $\tilde{\alpha}(\alpha):=\frac{\lceil m\alpha\rceil}{m}.$ Define the effective indices
$\alpha_{\mathrm{eff}}^{L}(\alpha)$ $:=\Big[\big(\bar{\gamma}_{L,m}-\varepsilon_{\mathrm{al}}(\alpha,\delta)\big)\tilde{\alpha}(\alpha)-b_{m}(\delta)\sqrt{\tilde{\alpha}(\alpha)}\Big]_{+},$
$\alpha_{\mathrm{eff}}^{U}(\alpha)$ $:=\Big[\big(\bar{\gamma}_{U,m}-\varepsilon_{\mathrm{al}}(\alpha,\delta)\big)\tilde{\alpha}(\alpha)-b_{m}(\delta)\sqrt{\tilde{\alpha}(\alpha)}\Big]_{+},$
where the remainder terms are defined as $\varepsilon_{\mathrm{al}}(\alpha,\delta):=\sqrt{\frac{\log(3m/\delta)}{2m\tilde{\alpha}(\alpha)}},$ and $b_{m}(\delta):=\sqrt{\frac{1}{2m}\log\frac{3m}{\delta}}.$ Then, with probability at least $1-\delta$ over $\mathcal{D}$, simultaneously for all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\alpha_{\mathrm{eff}}^{L}(\alpha)\big)\Big|\mathcal{D}\Big)$ $\geq 1-\alpha-\varepsilon_{m}(\delta),$
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\leq\hat{V}_{m}^{+}\big(1-\alpha_{\mathrm{eff}}^{U}(\alpha)\big)\Big|\mathcal{D}\Big)$ $\geq 1-\alpha-\varepsilon_{m}(\delta),$
where $\varepsilon_{m}(\delta):=\sqrt{\frac{\log(6/\delta)}{2m}}+\frac{1}{m}$. Let $V:[0,1]\to\mathbb{R}$ denote the quantile function of $\Delta(\psi)$. In particular, for any fixed $\tau\in(0,1)$, as $m\to\infty$, the band admits the heuristic form
$\hat{V}_{m}^{-}\big(\bar{\gamma}_{L,m}\tau\big)\lesssim V(\tau)\lesssim\hat{V}_{m}^{+}\big(\bar{\gamma}_{U,m}\tau+(1-\bar{\gamma}_{U,m})\big)$ (B.12)
B.4 Proof for Theorem B.1:
@@
We follow the setup of Theorem 3.1 and only provide the lower bound proof, since upper bound is exactly Theorem 3.1. Recall that we work conditionally on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j})$ and $\mathcal{G}:=\sigma\big(\{\mathcal{G}_{j}\}_{j=1}^{m}\big)$. For brevity, write $p_{j}:=p(\psi_{j})$, $q_{j}:=q(\psi_{j})$, and similarly $\hat{p}_{j},\hat{q}_{j}$. The true per-scenario discrepancy is $\Delta_{j}:=L(p_{j},\hat{q}_{j})$, $j=1,\dots,m$, and by the i.i.d. scenario assumption, $\{\Delta_{j}\}_{j=1}^{m}$ are i.i.d. draws from the distribution of $\Delta(\psi)$ under $\psi\sim\Psi$.
Define the lower pseudo-discrepancies
$\Delta_{j}^{-}:=\inf_{u\in\mathcal{C}_{j}(\hat{p}_{j})}L(u,\hat{q}_{j}),\qquad j=1,\dots,m,$
@@
where $\mathcal{C}_{j}(\hat{p}_{j})\subset\Theta$ satisfy the lower coverage guarantee $\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j})\mid\psi_{j},n_{j}\big)\geq\gamma_{L}.$ By Assumption 3.2 and compactness of $\mathcal{C}_{j}(\hat{p}_{j})$, Berge’s maximum theorem implies that the infimum is attained on $\mathcal{C}_{j}(\hat{p}_{j})$, hence $\Delta_{j}^{-}$ is well defined.
@@
Also by definition, on the event $\{p_{j}\in\mathcal{C}_{j}(\hat{p}_{j})\}$ we have $\Delta_{j}^{-}\leq\Delta_{j}$. Thus, for each $j$,
$\mathbb{P}\big(\Delta_{j}^{-}\leq\Delta_{j}\big|\mathcal{G}_{j}\big)\geq\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j})\big|\mathcal{G}_{j}\big)\geq\gamma_{L}.$
@@
Define the indicators $Y_{j}^{-}:=\mathbf{1}\{\Delta_{j}^{-}\leq\Delta_{j}\}$, $j=1,\dots,m$. Conditionally on $\mathcal{G}$, the variables $\{Y_{j}^{-}\}_{j=1}^{m}$ are independent and satisfy
$\mathbb{P}(Y_{j}^{-}=1\mid\mathcal{G})\geq\gamma_{L},\qquad\forall j.$ (B.13)
@@
We will fix $\alpha\in(0,1)$ for the moment, then generalize toward $\forall\alpha$. We again denote $\bar{V}_{m}(\alpha)$ as the empirical $\alpha$-quantile of the (unobservable) $\{\Delta_{j}\}_{j=1}^{m}$, i.e., $\bar{V}_{m}(\alpha):=\Delta_{(\lceil m\alpha\rceil)}$, where $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}$ are the order statistics, and let $\bar{F}_{m}(t):=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}\leq t\}$ be their empirical CDF. Define the index set of “lower-tail” scenarios
@@
By definition of $\bar{V}_{m}(\alpha)$ we have $s^{-}\geq\lceil m\alpha\rceil$. Among these, consider those for which the lower pseudo-gap is valid, where we denote $T_{\alpha}^{-}:=\{j\in S_{\alpha}^{-}:Y_{j}^{-}=1\}=\{j\in S_{\alpha}^{-}:\Delta_{j}^{-}\leq\Delta_{j}\leq\bar{V}_{m}(\alpha)\}.$ By (B.13),
$\mathbb{E}\big[|T_{\alpha}^{-}|\big|\mathcal{G}\big]=\sum_{j\in S_{\alpha}^{-}}\mathbb{P}(Y_{j}^{-}=1\mid\mathcal{G})\geq\gamma_{L}s^{-}.$
Applying Lemma A.2, for any $t\in[0,\gamma_{L}s^{-}]$,
$\mathbb{P}\Big(|T_{\alpha}^{-}|\leq\gamma_{L}s^{-}-t\Big|\mathcal{G}\Big)\leq\exp\Big(-\frac{2t^{2}}{s^{-}}\Big).$
@@
$\mathbb{P}\Big(|T_{\alpha}^{-}|\leq\gamma_{L}s^{-}-c\sqrt{s^{-}}\Big|\mathcal{G}\Big)\leq e^{-2c^{2}}.$ (B.14)
@@
$|U_{\alpha}^{-}|\geq|T_{\alpha}^{-}|\geq\gamma_{L}s^{-}-c\sqrt{s^{-}}.$
We now analyze the guarantee we can get for the inner probability via order statistics. Set $N:=\big\lfloor\gamma_{L}s^{-}-c\sqrt{s^{-}}\big\rfloor$ on the complement of the event in (B.14), so that $|U_{\alpha}^{-}|\geq N$ holds. Let $\hat{F}_{m}^{-}(t)$ denote the empirical CDF of $\{\Delta_{j}^{-}\}_{j=1}^{m}$, $\hat{F}_{m}^{-}(t):=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(\Delta_{j}^{-}\leq t),$ and define the empirical $\beta$-quantile $\hat{V}_{m}^{-}(\beta):=\inf\{t:\hat{F}_{m}^{-}(t)\geq\beta\}=\Delta^{-}_{(\lceil m\beta\rceil)}.$ If at least $N$ sample values satisfy $\Delta_{j}^{-}\leq\bar{V}_{m}(\alpha)$, then by Lemma A.7,
$\hat{V}_{m}^{-}(\beta)=\Delta^{-}_{(\lceil m\beta\rceil)}\leq\bar{V}_{m}(\alpha),$
whenever $\lceil m\beta\rceil\leq N$, i.e. whenever $\beta\leq N/m$.
A sufficient condition for $\beta\leq N/m$ is obtained by lower bounding $N/m$ using $s^{-}\geq\lceil m\alpha\rceil$. Indeed, on the same event,
$\frac{N}{m}\geq\frac{\gamma_{L}s^{-}-c\sqrt{s^{-}}-1}{m}\geq\gamma_{L}\alpha-c\sqrt{\frac{\alpha}{m}}-\frac{1}{m},$
where the last inequality uses $s^{-}\geq\lceil m\alpha\rceil\geq m\alpha$ and the floor/ceiling slack. Define
$\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m):=\gamma_{L}\alpha-c\sqrt{\frac{\alpha}{m}}-\frac{1}{m}.$ (B.15)
Then the preceding display implies $\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\leq N/m$, and hence
$\hat{V}_{m}^{-}\Big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\Big)\ \leq\ \bar{V}_{m}(\alpha)$ (B.16)
on the complement of the event in (B.14).
Thus, we define
$\mathcal{E}_{\alpha}^{-}$ $:=\Big\{\ \hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\ \leq\ \bar{V}_{m}(\alpha)\ \Big\},$ (B.17)
$\text{then}\qquad\mathbb{P}(\mathcal{E}_{\alpha}^{-})\ \geq\ 1-e^{-2c^{2}}.$
Next, we bridge $\bar{F}_{m}$ to the population CDF of $\Delta(\psi)$ under $\psi\sim\Psi$, which we denote by $F_{\Delta}$. By Lemma A.1, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
@@
where $\varepsilon_{m}(\delta):=\sqrt{\frac{\log(2/\delta)}{2m}}$.
Work on the intersection $\mathcal{E}_{\mathrm{DKW}}\cap\mathcal{E}_{\alpha}^{-}$. By (B.17) and the monotonicity of $F_{\Delta}$,
$F_{\Delta}\Big(\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big)\leq F_{\Delta}\big(\bar{V}_{m}(\alpha)\big).$
On $\mathcal{E}_{\mathrm{DKW}}$, we have
$F_{\Delta}\big(\bar{V}_{m}(\alpha)\big)\leq\bar{F}_{m}\big(\bar{V}_{m}(\alpha)\big)+\varepsilon_{m}(\delta)\leq\alpha+\varepsilon_{m}(\delta),$
where the last inequality uses $\bar{F}_{m}(\bar{V}_{m}(\alpha))\geq\alpha$ by definition of $\bar{V}_{m}(\alpha)$. Combining the last two displays yields
$F_{\Delta}\Big(\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big)\leq\alpha+\varepsilon_{m}(\delta).$
@@
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha-\varepsilon_{m}(\delta),$ (B.18)
with outer probability at least $1-\delta-e^{-2c^{2}}$.
We have shown that for any target level $\alpha$ and choice of $c>0$, the preceding argument yields a high-probability bound based on $\{\Delta_{j}^{-}\}_{j=1}^{m}$. We now extend the guarantee to hold uniformly over all $\alpha$. Fix $c>0$ and $\delta\in(0,1)$. For the grid $\alpha_{r}:=r/m$ ($r=1,\dots,m$), let
$\mathcal{E}_{\mathrm{DKW}}$ $:=\Big\{\sup_{t}\big|\bar{F}_{m}(t)-F_{\Delta}(t)\big|\leq\varepsilon_{m}(\delta)\Big\},$
@@
By DKW, $\Pr(\mathcal{E}_{\mathrm{DKW}})\geq 1-\delta$, and by the fixed-level argument, $\Pr(\mathcal{E}_{r}^{-})\geq 1-e^{-2c^{2}}$ for each $r$. Hence, by a union bound,